# LLM-DetectAIve: a Tool for Fine-Grained Machine-Generated Text Detection

### v1

@@ 5.1 Automatic Detection
The automatic detection interface is shown in Figure 1 (top). It allows users to input a text, and then the system responds with the class that the text belongs to. To ensure the prediction accuracy, the length of the submitted text is constrained to 50-500 words since the performance of our detectors drops significantly for shorter texts. Longer texts will be truncated by the tokenizer limited by the context window size of the BERT-like models.
@@ 5.2 Human Detector Playground
We developed the human detector playground as an interactive interface to engage users in the classification process. This feature allows users to test their own ability to distinguish between the four text categories. Figure 1 (bottom) shows the playground where users can explore the system, gaining insights into the subtle differences between various types of human and machine-generated texts.
@@ 5.3 Deployment and Implementation
The demo web application is deployed using Hugging Face Spaces. It is chosen for its seamless integration with transformer models, ease of use, and robust support for hosting machine learning applications. For implementing the user interface, we use Gradio. The code is publicly available and is licensed under the MIT License.
@@ 6 Conclusion and Future Work
In an era of advanced large language models, maintaining the integrity of text presents significant challenges. We presented a system that aims to identify the potential misuse of MGT, accurately differentiating human-written text from various types of machine text. Our system classifies text in a fine-grained manner — human-written, machine-generated, machine-polished, and machine-humanized texts, providing insights into the origins of the text and enhancing the trustworthy applications of LLMs. In future work, we plan to improve the Domain Adversarial Neural Network (DANN) to achieve more robust detection results. Additionally, we will explore the possibility of using a DANN on the text’s generator instead of the text’s domain to generalize detection across different text generators. Using a DANN on the domain and the generator could potentially lead to a truly universal detector. We further aim to add a fifth classification category: machine-written then human-edited text, to enhance our detection capabilities and to provide a more comprehensive analysis of text origins.
@@
We acknowledge certain limitations in this work that we plan to address in future research. First, although our work has explored more fine-grained machine-generated text scenarios beyond conventional binary classification, we did not consider a complex scenario where the text is first generated by machine and then manually edited by humans to suit their personal needs. This omission is primarily due to the high cost associated with collecting data that requires human editing. Additionally, issues were identified within the dataset. Specifically, some LLMs associate specific domains with particular formatting styles, such as markdown for lists, bullet points, and headers. This issue was particularly noticeable in the Wikihow and Peerread domains, where the LLMs frequently applied these formatting styles, potentially skewing the data and impacting the accuracy of our classifications. It also remains uncertain whether our system can generalize to detecting models or languages not included in our English only dataset.
@@
A primary ethical consideration is the data license. We reused pre-existing corpora, such as OUTFOX and Wikipedia, which have been publicly released and approved for research purposes. Furthermore, we generated new data on top of the original data, thereby mitigating concerns regarding data licensing.
@@
Considering that our data is generated by large language models, it might contain offensive or biased language.
@@
LLM-DetectAIve expands the conventional binary classification in MGT detection to more fine-grained levels, which is more aligned with real-life scenarios. We believe this approach could be applied in various scenarios, such as detecting students’ essays to ensure the originality of their work. Moreover, LLM usage detection may find applications in authorship detection in areas such as forensics.

### v3

@@ 5.1 Automatic Detection
The automatic detection interface is shown in Figure 1 (top). It allows users to input a text, and then the system responds with the class that the text belongs to. To ensure the prediction accuracy, the length of the submitted text is constrained to 50-500 words since the performance of our detectors drops significantly for shorter texts. Longer texts will be truncated, as we are limited by the context window size of the BERT-like transformers we use.
@@ 5.2 Human Detector Playground
The demo further offers a human detector playground as an interactive interface, which allows users to test their capability to distinguish between the four text categories. Figure 1 (bottom) shows a snapshot of the playground interface where the users can try the system, gaining insights into the subtle differences between various types of human-written and machine-generated texts.
@@ 5.3 Deployment and Implementation
Our demo is deployed on Hugging Face Spaces, which allows seamless integration with transformer models, ease of use, and robust support for hosting machine learning applications. For implementing the user interface, we used Gradio. The code is publicly available under an MIT license.
@@ 6 Conclusion and Future Work
In an era of advanced large language models, maintaining the integrity of text poses significant challenges. We presented a system that aims to identify the use of machine-generated text, accurately differentiating human-written text from various types of automatically generated text. Unlike previous work, we use a fine-grained classification schema (Human-Written, Machine-Generated, Machine-Written Machine-Humanized, and Human-Written Machine-Polished), which offers insights into the origins of the text, thus enabling trustworthiness. In future work, we plan to improve the Domain Adversarial Neural Network (DANN) to improve the results even further. We further plan to explore the possibility of using a DANN on the text’s generator instead of the text’s domain to generalize detection across different text generators. Using a DANN on both the domain and the generator could potentially lead to a truly universal detector. We also aim to expand the classification to include a fifth category: machine-written and human-edited text, enhancing detection capabilities and providing a more comprehensive analysis of text origins. To further improve the system, we also plan to address potential biases in the dataset caused by formatting styles linked to specific domains, such as Wikihow and PeerRead, to ensure better robustness across a broader range of human-written content. Last but not least, we want to expand the dataset to encompass a diverse set of languages, enabling the development of a robust multilingual detection model.
@@
We acknowledge certain limitations of our work, which we plan to address in future work. First, although our work has explored more fine-grained machine-generated text scenarios beyond conventional binary classification, we did not consider a complex scenario where the text is first generated by a machine and then is manually edited by humans to suit their personal needs. This is primarily due to the high costs associated with collecting data that requires human editing. Moreover, we identified some issues with the dataset. Specifically, some LLMs associate specific domains with particular formatting styles, such as markdown for lists, bullet points, and headers. This issue was particularly noticeable in the Wikihow and PeerRead domains, where the LLMs frequently applied these formatting styles, potentially skewing the data and impacting the accuracy of our classifications. It also remains uncertain whether our system can generalize to detecting models or languages not included in our English-only dataset.
@@
A primary ethical consideration is the data license. We reused pre-existing corpora, such as OUTFOX and Wikipedia, which have been publicly released and approved for research purposes. Moreover, we generated new data on top of the original data, thereby mitigating concerns regarding data licensing.
@@
Considering that our data is generated by large language models, it might contain offensive or biased language; we did not try to control for this, replying on the inetnal safety mechanisms of the LLMs we used.
@@
LLM-DetectAIve expands the conventional binary classification in machine-generated text detection to more fine-grained levels, which is more aligned with real-life scenarios. We believe this approach could be applied in various scenarios, e.g., for students’ essays to ensure the originality of their work. Moreover, LLM usage detection may find applications in authorship detection as well as in digital forensics.